# Grassroots Platforms with Atomic Transactions: Social Networks, Cryptocurrencies, and Democratic Federations

### v2

@@ 3. Grassroots Platforms with Atomic Transactions
Yet, the grassroots platforms we are interested in — grassroots social networks, grassroots cryptocurrencies, and grassroots federations — all call for atomic transactions for their specification. Therefore, we extended above the notion of a distributed transition system to include atomic transactions carried out by a set of agents. In this section we revisit the grassroots platforms previously defined using unary transition systems and redefined them more simply and abstractly using binary atomic transactions. In addition, we present a $k$-ary transactions-based specification of grassroots democratic federations, which where hitherto defined only at the community level.
3.1. Grassroots Social Network via Befriending and Defriending
A grassroots social network evolves by people forming and breaking friendships. The original definition of grassroots social networks (Shapiro, 2023b) was via a unary distributed transition system. Here both actions are specified—abstractly and concisely—as binary transactions. The local state of an agent $p\in P$ is the set of agents $Q\subseteq P$ of which $p$ is a friend.
@@ 3.1. Grassroots Social Network via Befriending and Defriending
We assume a given set of agents $P\subset\Pi$. The grassroots social network transition system over $P$ is the distributed transition system $SN$ over $P$, $S:=2^{P}$, and the binary transactions $c\rightarrow c^{\prime}$ over $\{p,q\}\subseteq P$ satisfying:
  1. (1)
@@
Proof.
The proof is by induction on the length of the run $r=c0,c_{1},\ldots,c_{n}$. Assume $|r|=1$, namely $r$ consists of the initial configuration $c0$. Then all local states are empty, satisfying the Lemma. Assume the lemma holds for runs of length $|r|=n$, that $c_{n}$ satisfies the Lemma, and consider the transition $c_{n}\rightarrow c_{n+1}$. It can be either Befriend or Defriend; both satisfy for every $p\in P$ the condition $p\in c_{q}$ iff $q\in c_{p}$ for $c_{n+1}$ if $c_{n}$ does. ∎
@@
In the unary implementation of these transactions, discussed in Section 5, befriending is realized by an offer by one person to another and a prompt response of accept or decline by the other person. Breaking a friendship is realized by a request by one of the two friends to the other and a prompt accept response by the other.
3.2. Grassroots Cryptocurrencies via Atomic Swaps
The original specification of grassroots cryptocurrencies was via unary distributed transition system and hence was quite involved, and so was the proof of them being grassroots (Shapiro, 2024). The definition in turn led to an implementation via a unary payment system (Lewis-Pye et al., 2023). Here we provide an alternative specification—abstract and concise—via binary transactions. We will later prove the specification to be grassroots.
@@ 3.2. Grassroots Cryptocurrencies via Atomic Swaps
We assume an infinite multiset $C$ of identical coins ¢. Given set of agents $P\subset\Pi$, we refer to ¢p, a $p$-indexed element of the indexed set $C^{P}$, as a $p$-coin, with ¢${}_{p}^{k}$ being a set of $k$ $p$-coins. A set of $P$-coins is a member of $2^{C^{P}}$, namely a multiset of $P$-indexed coins.
@@
Proof.
The proof is by induction on the length of the run $r=c0,c_{1},\ldots,c_{n}$. Assume $|r|=1$, namely $r$ consists of the initial configuration $c0$. Then all local states are empty, satisfying the Lemma. Assume the lemma holds for runs of length $|r|=n$, that $c_{n}$ satisfies the Lemma, and consider the transition $c_{n}\rightarrow c_{n+1}$. If the transition is Mint and the claim holds for all, then after minting it still holds by adding the minted coin. If it is Swap, it still holds as Swap does not adds new coins to the configuration or removes coins from it. ∎
3.3. Grassroots Federation via Joining and Leaving
Background. Grassroots federation (Halpern et al., 2024) is a process by which communities evolve and federate with other communities. Key steps include a community joining a federation, by the mutual consent of the two, and a community leaving a federation, by the unilateral decision of one of the two.
@@ 3.3. Grassroots Federation via Joining and Leaving
In a real deployment of grassroots federations we expect the partial order to be a superposition of geography, topics, and maybe other issues, so that the federation of dog lovers of a village can join both the general village federation and the regional federation of dog lovers, which in turn can join the general regional federation. The resulting federation structure, termed laminar, was investigated in (Halpern et al., 2024). With this constraint we can address the requirement that the federation graph always be acyclic.
@@
Naturally, this way many aspects of the transaction are abstracted-away, including that:
  1. (1)
@@
Definition 3.6 (Member, State, Projection).
Given a federation graph $G=(V,E\in\mathcal{G}(P)$, an agent $p\in P$ is a member of community $v\in G$ if there is a path in $G$ from $v$ to $p$. For a community $v\in V$, the state of $v$ in $G$ is the subgraph of $G$ that includes $v$, all edges in $G$ adjacent to $v$ and all nodes adjacent to these edges. For $p\in P$, let $G_{p}=(V_{p},E_{p})$, the projection of $G$ on $p$, denote the subgraph of $G$ that includes the state of all communities $p$ is a member of. Specifically, $V_{p}$ includes every node $v\in V$ with a directed path from $v$ to $p$ (namely all communities $p$ is a member of), as well as the state of $v$ in $G$. Also, for any $v\in V$, $P_{v}$ includes the members of $v$, namely $P_{v}:=\{p\in P:v\in V_{p}\}$.
Namely, $G_{p}$ includes any node adjacent to such a $v$ (namely all its children; its parents are already included due to having $p$ as a member), and $E_{p}$ is $E$ restricted to $V_{p}$.
Observation 3.7.
@@
Namely, the projections of a federation graph on its members fully define the graph. Hence, in the following grassroots federation transition system, the evolving federation graph $G$ created during a run is stored in a distributed way, with each agent $p\in P$ maintaining $G_{p}$. The result is that each agent stores the state of all communities they are a member of; equivalently, the state of each community is stored by all its members. Formally:
@@
This formulation can be considered as an abstraction of direct democracy, as every action of a community is carried out by all its members. This would mean that realizing this model requires all members of a community to jointly run a consensus protocol, the results of which define the acts of the community. While simple to specify, this approach may be unrealistic for large communities. For this and other reasons, works on grassroots federation (Halpern et al., 2024) explore a model in which each community is governed by a small (say, ${\approx}100$) assembly, selected by sortition among the members of the community. In such a model, only assembly members would store the state of the community, and only the assembly, not the entire community, would run the consensus protocol realizing its behaviour. It is important to note that in both formulations, the consensus (between the entire population, or just the assembly) is on the decision taken. In case of an assembly, the decision process itself could be “on-chain”, e.g. realized by a social contract realizing the various methods developed for the egalitarian democratic formation and governance of digital communities (Poupko et al., 2021; Shahaf et al., 2020, 2018; Meir et al., 2024; Abramowitz et al., 2021b; Bulteau et al., 2021; Elkind et al., 2021; Abramowitz et al., 2021a), or “of-chain”, e.g. via an online video meeting of the members of the assembly in which votes are taken by show of hands. The only requirement is that there will be consensus among the members of the assembly on the democratic decision that was taken. Here, we stick with the direct-democracy model to simplify the exposition and are agnostic/abstract-away the decision process employed.
@@
Examining these changes shows that they all preserve validity, in that also in the new configuration each agent $p$ records, by construction, exactly the $p$-projection of the updated federation graph $G^{\prime}$. ∎
This completes the transactions-based specification of the three key grassroots platforms. It does not distinguish between transactions carried out by mutual consent (befriend, open a credit line, join), and transactions that are obligatory once requested (defriend, redeem, leave). The distinction can be added at this level of abstraction, with a liveness requirement that once a request to carry out an obligatory transaction is made, it is eventually carried out. Alternatively, it can be added at the unary implementation level; this may be more natural as atomic transactions would anyway be realised by one agent issuing a request/offer and the other(s) consenting to it (Shapiro, 2023b; Lewis-Pye et al., 2023).
4. Grassroots Protocols
@@
4.1. Protocols and Grassroots Protocols
@@
Note that $c_{p}$, the state of $p$ in $c$, is in $S(P^{\prime})$, not in $S(P)$, and hence may include elements “alien” to $P$, e.g., friendship with or a coin of $q\in P^{\prime}\setminus P$.
@@
A protocol $\mathcal{F}$ is: (1) oblivious if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$, $T(P){\uparrow}P^{\prime}\subseteq T(P^{\prime})$ (2) interactive if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$ and every configuration $c\in C(P^{\prime})$ such that $c{/}P\in C(P)$, there is a computation $c\xrightarrow{*}c^{\prime}$ of $\mathcal{F}(P^{\prime})$ for which $c^{\prime}{/}P\notin C(P)$. (3) grassroots if it is oblivious and interactive.
@@
Being oblivious implies that if a run of $\mathcal{F}(P^{\prime})$ reaches some configuration $c$, then anything $P$ could do on their own in the configuration $c/P$ (with a transition from $T(P))$, they can still do in the larger configuration $c$ (with a transition from $T(P^{\prime})$). The reason is that if $t=(c/P\rightarrow c^{\prime})\in T(P)$, then by the definition of closure, $t^{\prime}=(c\rightarrow c^{\prime\prime})\in t{\uparrow}P^{\prime}\subseteq T(P^{\prime})$, where $c^{\prime\prime}_{p}=c^{\prime}_{p}$ if $p\in P$ else $c^{\prime\prime}_{p}=c_{p}$. Inductively, this implies that if agents $P$ employ only transitions in $T(P)$ from the start, with their local states remaining in $S(P)$, they could continue doing so indefinitely, effectively ignoring any members in $P^{\prime}\setminus P$. Proving the corresponding claim within the original definition was more difficult, and required the notion of interleaving of computations of two distributed transition systems (Shapiro, 2021).
@@
We note that any protocol that uses a global data structure, whether replicated (Blockchain) or distributed (DHT (Rhea et al., 2005), IPFS (Benet, 2014)), is not oblivious as members of $P$ cannot ignore changes to the global data structure made by members of $P^{\prime}\setminus P$; and hence is also not grassroots.
@@
A slight complication in the definition of interactive is the use of a finite computation $\xrightarrow{*}$ rather than just a single transition $\rightarrow$. It is required as in some platforms agents need to make some preparatory steps before they may interact with each other. For example, in grassroots cryptocurrencies agents need first to mint some coins before they can swap them.
@@
Being interactive is a weak liveness requirement. Informally, the standard liveness requirement is “something good must eventually happen”. Here, the requirement is that “it must be the case that something good ($P$ interacting with non-$P$) may eventually happen”. Namely, no matter what members of $P$ do, it is always the case that they can eventually interact with non-$P$’s. Moreover, being interactive not only requires that $P$ can always eventually interact with $P^{\prime}\setminus P$, but that they do so in a way that leaves “alien traces” in the local states of $P$, so that the resulting configuration $c^{\prime}/P$ could not have been produced by $P$ running on their own. For example, forming friendships with, or receiving coins from, agents outside of $P$.
4.2. Transactions-Based Grassroots Protocols
@@
The original paper (Shapiro, 2023a) provided sufficient conditions for a protocol to be grassroots—asynchrony, interference-freedom, and interactivity. They should still hold under the new definition, with some adaptation, mostly simplification. However, here we are interested in transactions-based transition systems, therefore we will follow a different route: We first show how a local-states function can be used to define a set of transactions, and then show how a set of such transactions can be used to define a protocol, termed transactions-based protocol. We then prove that a single condition on such transactions—interactivity—is sufficient for a transactions-based protocol to be grassroots.
@@
Next we aim to find conditions under which a transactions-based protocol is grassroots. In the original paper (Shapiro, 2023a), fulfilling three conditions were deemed sufficient: Asynchrony, interference-freedom, and interactivity. Intuitively speaking, transactions-based distributed transition systems are asynchronous by construction, as the essence of a transaction is that it can be taken no matter what the states of non-participants are. They are also non-interfering for the same reason. The following Proposition captures this:
@@
by definition of $T(P)$, Lemma 4.7, and since $P\subset P^{\prime}$ and therefore $R(P)\subseteq R(P^{\prime})$. ∎
@@
Proof.
Let $S$ be a local-states function, $R$ an interactive set of transactions over $S$, and $\mathcal{F}$ a transactions-based protocol over $R$ and $S$, $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$, and $c\in C(P^{\prime})$ a configuration such that $c{/}P\in C(P)$. By $R$ being interactive (Definition 4.8), there is a computation $c\xrightarrow{*}c^{\prime}\subseteq R(P^{\prime}){\uparrow}P^{\prime}$ for which $c^{\prime}{/}P\notin C(P)$. By the definition of $\mathcal{F}$ as a transactions-based protocol over $R$ and $S$, $T(P^{\prime})=R(P^{\prime}){\uparrow}P^{\prime}$, hence this computation is of $T(P^{\prime})$, which is thus interactive. ∎
@@
Proof.
Let $\mathcal{F}$ be a protocol over a set of transactions $R$ (Definition 4.5), where $R$ is interactive (Definition 4.8). Since $\mathcal{F}$ is a transactions-based protocol then, according to Proposition 4.6, $\mathcal{F}$ is oblivious. And since $\mathcal{F}$ is over an interactive set of transactions then, according to Observation 4.9, $\mathcal{F}$ is interactive. Therefore, by Definition 4.3, $\mathcal{F}$ is grassroots.. ∎
4.3. The Three Platforms are Grassroots as Specified

### v3

@@ 3 Grassroots Platforms with Atomic Transactions
Yet, the grassroots platforms we are interested in—grassroots social networks, grassroots cryptocurrencies, and grassroots democratic federations—all call for atomic transactions for their specification. Therefore, we extended above the notion of a distributed transition system to include atomic transactions carried out by a set of agents. In this section we revisit the grassroots platforms previously defined using unary transition systems and redefine them more simply and abstractly using binary atomic transactions. In addition, we present a $k$-ary transactions-based specification of grassroots democratic federations, which where hitherto defined only at the community level.
3.1 Grassroots Social Network via Befriending and Defriending
A grassroots social network evolves by people forming and breaking friendships. The original definition of grassroots social networks [44] was via a unary distributed transition system. Here both actions are specified—abstractly and concisely—as binary transactions. The local state of an agent $p\in P$ is the set of agents $Q\subseteq P$ of which $p$ is a friend.
@@ 3.1 Grassroots Social Network via Befriending and Defriending
We assume a given set of agents $P\subset\Pi$. The grassroots social network transition system over $P$ is the distributed transition system $SN$ over $P$, $S:=2^{P}$, and the binary transactions $c\rightarrow c^{\prime}$ with every $\{p,q\}\subseteq P$ as participants, satisfying:
  1. 1.
@@
Proof 3.3.
The proof is by induction on the length of the run $r=c0,c_{1},\ldots,c_{n}$. Assume $|r|=1$, namely $r$ consists of the initial configuration $c0$. Then all local states are empty, satisfying the Lemma. Assume the lemma holds for runs of length $|r|=n$, that $c_{n}$ satisfies the Lemma, and consider the transition $c_{n}\rightarrow c_{n+1}$. It can be either Befriend or Defriend; both satisfy for every $p\in P$ the condition $p\in c_{q}$ iff $q\in c_{p}$ for $c_{n+1}$ if $c_{n}$ does.
@@
3.2 Grassroots Cryptocurrencies via Atomic Swaps
The original specification of grassroots cryptocurrencies was via unary distributed transition system and hence was quite involved, and so was the proof of them being grassroots [45]. The definition in turn led to an implementation via a unary payment system [29]. Here we provide an alternative specification—abstract and concise—via binary transactions. We will later prove the specification to be grassroots.
@@ 3.2 Grassroots Cryptocurrencies via Atomic Swaps
We assume an infinite multiset $C$ of identical coins ¢. Given set of agents $P\subset\Pi$, we refer to ¢p, a $p$-indexed element of the indexed set $C^{P}$, as a $p$-coin, with ¢${}_{p}^{k}$ being a multiset of $k$ $p$-coins. A set of $P$-coins is a member of $2^{C^{P}}$, namely a multiset of $P$-indexed coins.
@@
Proof 3.7.
The proof is by induction on the length of the run $r=c0,c_{1},\ldots,c_{n}$. Assume $|r|=1$, namely $r$ consists of the initial configuration $c0$. Then all local states are empty, satisfying the Lemma. Assume the lemma holds for runs of length $|r|=n$, that $c_{n}$ satisfies the Lemma, and consider the transition $c_{n}\rightarrow c_{n+1}$. If the transition is Mint and the claim holds for all, then after minting it still holds by adding the minted coin. If it is Swap, it still holds as Swap does not adds new coins to the configuration or removes coins from it.
3.3 Grassroots Democratic Federation via Joining and Leaving
Background. Grassroots federation [25] is a process by which communities evolve and federate with other communities. Key steps include a community joining a federation, by the mutual consent of the two, and a community leaving a federation, by the unilateral decision of one of the two.
@@ 3.3 Grassroots Democratic Federation via Joining and Leaving
In a real deployment of grassroots federations we expect the partial order to be a superposition of geography, topics, and maybe other issues, so that the federation of dog lovers of a village can join both the general village federation and the regional federation of dog lovers, which in turn can join the general regional federation. The resulting federation structure, termed laminar, was investigated in [25]. With this constraint we can address the requirement that the federation graph always be acyclic.
@@
Naturally, this results in many aspects of the transaction being abstracted-away, including that:
  1. 1.
@@
Definition 3.8 (Community and Personal Subgraph).
Given a federation graph $G=(V,E\in\mathcal{G}(P)$, For a community $v\in V$, the community subgraph $G_{v}=(V_{v},E_{v})$ of $v$ is the subgraph of $G$ where $V_{v}$ includes $v$ and all nodes adjacent to $v$ and $E_{v}$ includes all edges incident with $v$. An agent $p\in P$ is a member of community $v\in G$ if there is a path in $G$ from $v$ to $p$, and the personal subgraph $G_{p}$ of $p$ is the union of the subgraphs of the communities $p$ is a member of, namely $G_{p}:=\bigcup_{p\in v}G_{v}$
Thus, $G_{p}$ includes any community $v$ that $p$ is a member of and any edge incident with $v$
{observation}
@@
Namely, a federation graph is fully-defined by its members’ subgraphs. Hence, in the following grassroots federation transition system, the evolving federation graph $G$ created during a run is stored in a distributed way, with each agent $p\in P$ maintaining $G_{p}$. The result is that each agent stores the state of all communities they are a member of; equivalently, the state of each community is stored by all its members. Formally:
@@
This formulation can be considered as an abstraction of direct democracy, as every action of a community is carried out by all its members. This would mean that realizing this model requires all members of a community to jointly run a consensus protocol, the results of which define the acts of the community. While simple to specify, this approach may be unrealistic for large communities. For this and other reasons, works on grassroots federation [25] explore a model in which each community is governed by a small (say, ${\approx}100$) assembly, selected by sortition among the members of the community. In such a model, only assembly members would store the state of the community, and only the assembly, not the entire community, would run the consensus protocol realizing its behaviour. It is important to note that in both formulations, the consensus (between the entire population, or just the assembly) is on the decision taken. In case of an assembly, the decision process itself could be “on-chain”, e.g. implemented by a social contract [13] realizing the various methods developed for the egalitarian democratic formation and governance of digital communities [37, 41, 40, 33, 3, 11, 22, 2], or “of-chain”, e.g. via an online video meeting of the members of the assembly in which votes are taken by show of hands. The only requirement is that there will be consensus among the members of the assembly on the democratic decision that was taken. Here, we stick with the direct-democracy model to simplify the exposition and are agnostic/abstract-away the decision process employed.
@@
Examining these changes shows that they all preserve validity, in that also in the new configuration each agent $p$ records, by construction, exactly the $p$-projection of the updated federation graph $G^{\prime}$.
This completes the transactions-based specification of the three key grassroots platforms. It does not distinguish between transactions carried out by mutual consent (befriend, open a credit line, join), and transactions that are obligatory once requested (defriend, redeem, leave). The distinction can be added at this level of abstraction, with a liveness requirement that once a request to carry out an obligatory transaction is made, it is eventually carried out. Alternatively, it can be added at the unary implementation level; this may be more natural as atomic transactions would anyway be realised by one agent issuing a request/offer and the other(s) consenting to it [44, 29].
4 Grassroots Protocols
@@
4.1 Protocols and Grassroots Protocols
@@
Note that in the definition above, $c_{p}$, the state of $p$ in $c$, is in $S(P^{\prime})$, not in $S(P)$, and hence may include elements “alien” to $P$, e.g., friendship with or a coin of $q\in P^{\prime}\setminus P$.
@@
A protocol $\mathcal{F}$ is: 1. oblivious if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$, $T(P){\uparrow}P^{\prime}\subseteq T(P^{\prime})$ 2. interactive if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$ and every configuration $c\in C(P^{\prime})$ such that $c{/}P\in C(P)$, there is a computation $c\xrightarrow{*}c^{\prime}$ of $\mathcal{F}(P^{\prime})$ for which $c^{\prime}{/}P\notin C(P)$. 3. grassroots if it is oblivious and interactive.
@@
Oblivious. Being oblivious implies that if a run of $\mathcal{F}(P^{\prime})$ reaches some configuration $c^{\prime}$, then anything $P$ could do on their own in the configuration $c^{\prime}/P$ (with a transition from $T(P))$, they can still do in the larger configuration $c^{\prime}$ (with a transition from $T(P^{\prime})$), effectively being oblivious to members of $P^{\prime}\setminus P$. The reason is that if $t=(c^{\prime}/P\rightarrow d)\in T(P)$, then by the definition of closure, $t^{\prime}=(c^{\prime}\rightarrow d^{\prime})\in t{\uparrow}P^{\prime}\subseteq T(P^{\prime})$, where $d^{\prime}_{p}=d_{p}$ if $p\in P$ else $d^{\prime}_{p}=c^{\prime}_{p}$. Inductively, this implies that if agents $P$ employ only transitions in $T(P)$ from the start, with their local states remaining in $S(P)$, they could continue doing so indefinitely, effectively ignoring any members in $P^{\prime}\setminus P$. Proving the corresponding claim within the original definition was more difficult, and required the notion of interleaving of computations of two distributed transition systems [42].
@@
We note that any protocol that uses a global data structure, whether replicated (Blockchain) or distributed (DHT [39], IPFS [6]), is not oblivious as members of $P$ cannot ignore changes to the global data structure made by members of $P^{\prime}\setminus P$; and hence is not grassroots.
@@
Interactive. A slight complication in the definition of interactive is the use of a finite computation $\xrightarrow{*}$ rather than just a single transition $\rightarrow$. It is required as in some grassroots platforms agents need to make some preparatory steps before they may interact with each other. For example, in grassroots cryptocurrencies agents need first to mint some coins before they can swap them.
@@
Being interactive is a weak liveness requirement. Informally, a standard liveness requirement has the form “something good must eventually happen”. Here, the requirement is “it must be the case that something good ($P$ interacting with non-$P$) may eventually happen”. Namely, no matter what members of $P$ do, it is always the case that they can eventually interact with non-$P$’s. Moreover, being interactive not only requires that $P$ can always eventually interact with $P^{\prime}\setminus P$, but that they do so in a way that leaves “alien traces” in the local states of $P$, so that the resulting configuration $c^{\prime}/P$ could not have been produced by $P$ running on their own. For example, forming friendships with, or receiving coins from, agents outside of $P$. We note that while federated systems [38, 7] are oblivious, as one server may choose to ignore all other servers and just serve its clients, they are not interactive, as a set of clients $P$ without a server cannot do more if embedded within a larger set of clients $P^{\prime}$, still without a server.
4.2 Transactions-Based Grassroots Protocols
@@
The original paper [43] provided sufficient conditions for a protocol to be grassroots—asynchrony, interference-freedom, and interactivity. They should still hold under the new definition, with some adaptation, mostly simplification. However, here we are interested in transactions-based transition systems, therefore we will follow a different route: We first show how a local-states function can be used to define a set of transactions, and then show how a set of such transactions can be used to define a protocol, termed transactions-based protocol. We then prove that a single condition on such transactions—interactivity—is sufficient for a transactions-based protocol to be grassroots.
@@
Next we aim to find conditions under which a transactions-based protocol is grassroots. In the original paper [43], fulfilling three conditions were deemed sufficient: Asynchrony, interference-freedom, and interactivity. Intuitively speaking, transactions-based distributed transition systems are asynchronous by construction, as the essence of a transaction is that it can be taken no matter what the states of non-participants are. They are also non-interfering for the same reason. The following Proposition captures this:
@@
by definition of $T(P)$, Lemma 4.7, and since $P\subset P^{\prime}$ and therefore $R(P)\subseteq R(P^{\prime})$.
@@
Proof 4.12.
Let $S$ be a local-states function, $R$ an interactive set of transactions over $S$, and $\mathcal{F}$ a transactions-based protocol over $R$ and $S$, $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$, and $c\in C(P^{\prime})$ a configuration such that $c{/}P\in C(P)$. By $R$ being interactive (Definition 4.10), there is a computation $c\xrightarrow{*}c^{\prime}\subseteq R(P^{\prime}){\uparrow}P^{\prime}$ for which $c^{\prime}{/}P\notin C(P)$. By the definition of $\mathcal{F}$ as a transactions-based protocol over $R$ and $S$, $T(P^{\prime})=R(P^{\prime}){\uparrow}P^{\prime}$, hence this computation is of $T(P^{\prime})$, establishing that $\mathcal{F}$ is interactive.
@@
Proof 4.14.
Let $\mathcal{F}$ be a protocol over a set of transactions $R$ (Definition 4.5), where $R$ is interactive (Definition 4.10). Since $\mathcal{F}$ is a transactions-based protocol then, according to Proposition 4.6, $\mathcal{F}$ is oblivious. And since $\mathcal{F}$ is over an interactive set of transactions then, according to Proposition 4.11, $\mathcal{F}$ is interactive. Therefore, by Definition 4.3, $\mathcal{F}$ is grassroots..
4.3 The Three Platforms are Grassroots as Specified